OpenStudy (amistre64):

Why is a broken equation editor a new feature? Seems to me to be a step backwards; gonna have to start piling rocks together to express numerals and work my way thru roman numberals ... ;)

OpenStudy (amistre64):

attach file seems to work still; but no equation editor....

OpenStudy (anonymous):

._. LOL , I like the rock idea (more of a manual person) ^_^ :D

OpenStudy (anonymous):

but still, didn't get what the attachment meant .-.

OpenStudy (amistre64):

the attachment means that the attach file button still works; but the equation button is broken

OpenStudy (amistre64):

and it was the only pic i could find on the schools computer

OpenStudy (anonymous):

LOL, true, so I wasn't the only one facing such problem

OpenStudy (amistre64):

i think there is a conspiricay afoot to limit our ability to answer quaestions in a way that resembles a modicum of intelligence lol

OpenStudy (shadowfiend):

Hmm... What browser?

OpenStudy (shadowfiend):

Also, what's broken about it?

OpenStudy (anonymous):

I think it still works. Can't figure out how to write fractions.

OpenStudy (shadowfiend):

(I just tried it and it seems to work fine for me, so want to get a bit more info as I try and find the issue.)

OpenStudy (shadowfiend):

chaguanas -- yeah, I still need to get a toolbar item for fractions in there. You can do fractions in two ways. If the fraction is over the entire thing, you can use \over. e.g.: 2x^2 + 3x + 2 \over 3x^2 + 2x - 9 is: \[2x^2 + 3x + 2 \over 3x^2 + 2x - 9\] If you have multiple equations, you can use \frac{numerator}{denominator}, e.g.: \frac{1}{2}x^2 + 3x + \frac{2}{x^3} is: \[\frac{1}{2}x^2 + 3x + \frac{2}{x^3}\]

OpenStudy (anonymous):

\[oh\over good\]

OpenStudy (shadowfiend):

Ok, I just found the issue with the math -- looks like it's IE only. Will try and get it fixed for tomorrow's push; sorry about that!

OpenStudy (amistre64):

firefox aint doin nothing for me... it doesnt even translate it from the textbox...

OpenStudy (amistre64):

\[\sqrt{3+1}\]

OpenStudy (amistre64):

nuthin...

OpenStudy (shadowfiend):

Hm. K. (Hehe, mine renders it perfectly :p)

OpenStudy (shadowfiend):

Yeah, I think I know why this is happening.

OpenStudy (amistre64):

knowing and fixing .... 2 different things lol. Its still broke on my end, and others as well...

OpenStudy (anonymous):

yeah..it isnt working at my end too

OpenStudy (watchmath):

test \[a^2+b^2\] too heavy for mathjax shadowfiend?

OpenStudy (watchmath):

No LaTeX no fun :( (I am going off) How if you use image instead shadowfiend? for example using mimetex.cgi. It is uglier than mathjax but it is lightweight.

OpenStudy (watchmath):

Ok, it works in firefox but not in Chrome :).

OpenStudy (amistre64):

it aint working in firefox....

OpenStudy (shadowfiend):

Oh, sorry -- when I said I knew why it was happening I didn't mean I was going to fix it immediately :) Everything should be working now. It's interesting because it seems like which browsers broke was inconsistent. Anyway, everything's fixed up. We switched to a new version of MathJax and their way of loading the configuration file changed, but I failed to change along with it. I'm guessing in certain situations it went ahead and loaded the default configuration, and in others it failed to.

OpenStudy (watchmath):

So are you using MathJax 1.1 ?

OpenStudy (shadowfiend):

Yep. There are still some bugs with rendering in real time, unfortunately, which I'll be pinging the MathJax folks about.

OpenStudy (shadowfiend):

And sorry, to respond to your image suggestion: images are more load on the server, and generally look worse. The upside is, they're a bit less buggy. But, we're huge fans of MathJax. Wouldn't give it up for the world :)